OpenStudy (anonymous):

Which of the following equations could be the result of using the comparison method to solve the system shown? x + 2y = 6 x - 4y = 8 2y - 6 = 4y + 8 6 - 2y = -4y + 8 6 - 2y = 4y + 8 i got a

OpenStudy (anonymous):

it's the third one. x + 2y = 6 <=> x = 6 - 2y x - 4y = 8 <=> x = 8 + 4y so 6-2y = 8+4y